Then diago- 6. Find the eigenvalues and eigenvectors of the matrix A = nalize the matrix. [4 points)
Answer #1

Ane) A = [1 s find eign value The cign value are those a for which IA-XIO *** det (a-at) - det([15] -66:]) det([*?3-0 : so -Find eign Vector if IA-al=0 then the equation b=0 (A-AI) x = b AX AX for A=3 the oan AX=ax becomes AX=3x writing x=(x 7 AX= [entrere t es any voal No..the nonzero vectory that Sabesfy AX = -3x que called eign vector associated with the eign value a=3Diagonalize the matrix A=[-2-5 by PDP-1=A the eignvalue of A are 34-1. correspond eign vector are alle Esmosbectivity, P-E-5]